Mark volunteers a local pet store. He made took some notes about the animals in the store one week. 2/5 are birds 3/8 are bunnies. The rest are other pets, such as fish, snakes,and mice.
What fraction of the animals in the store are birds and bunnies.
What fraction of the in the store are not birds and bunnies.

Answer by greenestamps(13198)



What fraction of the animals in the store are birds and bunnies?

There are no animals in the store that are both birds and bunnies; so the answer is 0.

What fraction of the animals in the store are not birds and bunnies?

ALL of them are not both birds and bunnies; so the answer is 1.

Now let's ask and answer the right questions.

What fraction of the animals in the store are birds OR bunnies?

birds 2/5; bunnies 3/8. Get a common denominator: 5*8 = 40.

birds 16/40; bunnies 15/40

birds OR bunnies: 16/40+15/40 = 31/40

What fraction of the animals in the store are not birds OR bunnies?

1-31/40 = 9/40